5) a) What is the present value of S40 earned 2-years from now if compounding was semi-annual and the interest rate is annually 3%? b) A black boxjust paid $20, which is expected to grow by 3% when the interest rate is 7% forever, what is the present value of this black box? c) What is the future value of an annuity due with a $15 cash flow, 4% annual interest with quarterly compounding three-years from now? d) If the monthly interest rate is 0.5%, then what is the effective annual rate? What is the stated annual rate?
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins as Cut ta copy. Σ AutoSum ー E ゴWrap Text в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Sort &Find & Format Painter Formatting as Table Styles2 Clear Clipboard DS440 DM Alignment Number Cells Edting DN DO DP DQ DR DS DT DU DV DW 432 433 434 435 436 437 438 439 440 441 442 443 1 FV 40 YEARS RATE PERIODS IN A YEAR 396 SEMI-ANNUAL 37.69 [EXCEL PV(DQ434/DQ435,2 DQ435,,-DQ432)] 2 IT IS A CASE OF PRESENT OF GROWING PERPETUITY PVGA 20(1+0.03)/(0.07-0.03)-515 A AMOUNT RECEIVED AFTER A YEAR PVGA A/(r-g) PV= 20 + PVGA-20 + 515 = 535 3 FUTURE VALUE OF ANNUITY DUE PMT 15 RATE EAR =(1+0.04/4)^4-1 FVAD = 4% QUARTERLY COMPOUNDING (m-4) 4.05% (1+/m)^m-1 445 446 447 448 449 450 11 1 トーCALCULATOR ( LOAN OPTIONS | pv, fv, annuity i CAP STRU VALUE $48.75 EXCEL-FV(DQ445,DQ443,-DQ442,1)] 0.50% m-12, r/m =0.5% 6.17% (1+r/m)^m-1 6.00% (0.5%*12) 4 MONTHLY RATE EAR-(1+0.005)A12-1 STATED ANNUAL RATE BOX DILUTION DOLLAR COST AVGMORTGAGE EXPO Sheeti Sheet2 ShelI rences: x261 02:59